# Inv-Entropy: A Fully Probabilistic Framework for Uncertainty Quantification in Language Models

### v1

@@ Abstract
Large language models (LLMs) have transformed natural language processing, but their reliable deployment requires effective uncertainty quantification (UQ). Existing UQ methods are often heuristic and lack a probabilistic foundation. This paper begins by providing a theoretical justification for the role of perturbations in UQ for LLMs. We then introduce a dual random walk perspective, modeling input–output pairs as two Markov chains with transition probabilities defined by semantic similarity. Building on this, we propose a fully probabilistic framework based on an inverse model, which quantifies uncertainty by evaluating the diversity of the input space conditioned on a given output through systematic perturbations. Within this framework, we define a new uncertainty measure, Inv-Entropy. A key strength of our framework is its flexibility: it supports various definitions of uncertainty measures, embeddings, perturbation strategies, and similarity metrics. We also propose GAAP, a perturbation algorithm based on genetic algorithms, which enhances the diversity of sampled inputs. In addition, we introduce a new evaluation metric, Temperature Sensitivity of Uncertainty (TSU), which directly assesses uncertainty without relying on correctness as a proxy. Extensive experiments demonstrate that Inv-Entropy outperforms existing semantic UQ methods. The code to reproduce the results can be found at https://github.com/UMDataScienceLab/Uncertainty-Quantification-for-LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success in various natural language processing tasks, such as text generation, question answering, and summarization (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023). These models, particularly driven by recent advances in transformers and attention mechanisms, have pushed the boundaries of what is achievable in language understanding and generation (Chung et al., 2024; OpenAI, 2023). However, despite their impressive capabilities, a significant challenge remains: LLMs tend to hallucinate, or generate confidently wrong predictions (Maynez et al., 2020; Zhang et al., 2024b). This is a serious concern in applications where reliability is paramount, such as healthcare, autonomous systems, and legal domains, where incorrect outputs can have dire consequences (Ji et al., 2023). Addressing these challenges is essential for ensuring the reliable and responsible deployment of LLMs at scale, ultimately unlocking their full potential in real-world applications.
@@
A central step in addressing these limitations is developing effective measures for UQ, enabling LLMs to acknowledge their confidence in a generated output. Existing UQ approaches rely predominantly on heuristic consistency checks. Most commonly, they use the model’s own generation likelihood or perplexity as a proxy for confidence, or they measure dispersion across multiple sampled continuations (e.g. via n-gram overlap or embedding-space variance) (Mudumbai and Bell, 2024). Such likelihood and sampling-based measures, however, lack a grounded probabilistic interpretation. Also, token-level probabilities are known to dramatically under-estimate uncertainty as models can be “confidently wrong” (Jiang et al., 2021), and they are often inaccessible in black-box LLMs.
@@
To probe deeper model brittleness, recent work has turned to input-perturbation UQ methods. Variant prompts are created through paraphrasing, adversarial token insertion, or temperature shifts. Output sensitivity is then quantified as an uncertainty signal (Gao et al., 2024; Tuna et al., 2022; Seeböck et al., 2019). For instance, Gao et al. (2024) randomly perturb both prompt wording and sampling temperature, then aggregate variation across outputs to flag unstable predictions. Tuna et al. (2022) applies adversarial paraphrases to uncover “blind spots” where small semantic-preserving edits trigger large output changes, while Seeböck et al. (2019) uses systematic character-level and word-level corruptions to map regions of high model vulnerability.
@@
Despite these advances, a fundamental probabilistic gap remains, as existing methods mainly rely on heuristic score functions for UQ. To address this, we introduce Inv-Entropy, a fully probabilistic framework grounded in random-walk theory that offers a new perspective: it learns the distribution of possible inputs that could have produced a given outcome. This is accomplished through structured perturbations, which simultaneously capture input variability and the influence of different inputs on model predictions. Our contributions are summarized as follows:
@@
    We present the first work in UQ for LLMs that adopts a fully probabilistic framework, grounded in random walk theory. This framework is highly flexible, and its probabilistic nature allows for the use of various UQ measures. It is also intrinsically capable of handling black-box models, as it is agnostic to token probabilities.
@@
    Theoretically, our work provides a rigorous foundation for using perturbation-based methods for UQ, validating the recent momentum behind such approaches.
@@ 2.1 Why perturb the input?
Unlike model-free UQ approaches (Shafer and Vovk, 2008), we exploit a key property specific to language: semantic equivalence. An ideal LLM should respond similarly to semantically equivalent inputs. Therefore, we can design an approach based on input perturbations on the semantic equivalent class to detect inconsistencies in LLM outcomes.
Figure 1: Left: level sets of the ground truth $f^{\star}$. We perturb input $x_{0}$ to $x_{1},x_{2},\cdots$ on the isocontour so that $f^{\star}(x_{0})=f^{\star}(x_{1})=\cdots$. Right: level sets of the model $\hat{f}$. The deviations from $\hat{f}(x_{i})_{i\geq 1}$ to $\hat{f}(x_{0})$ characterize the uncertainty of $\hat{f}$.
To highlight the importance of input perturbation, we start from a proof-of-concept example where the target function $f^{\star}$ is a single output function $f^{\star}:\mathbb{R}^{d}\to\mathbb{R}$. We don’t have access to $f^{\star}$, but we have access to an approximation $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}$, such as a pre-trained model. At first sight, it seems hopeless to quantify the alignment between $\hat{f}$ and $f^{\star}$ without knowing $f^{\star}$. However, the semantic equivalence class relative to an input $x_{0}$ provides a set $\mathcal{I}(x_{0})\subseteq\mathbb{R}^{d}$, such that $f^{\star}(x^{\prime})=f^{\star}(x_{0}),\,\forall x^{\prime}\in\mathcal{I}(x_{0})$. The equivalence class provides valuable information for UQ. An illustration is shown in Fig. 1.
The shape of $\mathcal{I}(x_{0})$ could be complex for general $f^{\star}$. However, in a small neighborhood of $x_{0}$ where $\nabla f^{\star}(x_{0})\neq 0$, $\mathcal{I}(x_{0})$ should be locally close to the tangent space of $f^{\star}$. More specifically, we define tangent invariance set as $\mathcal{I}_{\text{tangent}}(x_{0})=\{x_{0}+\left(I-\eta_{\nabla f^{\star}}\right)z;z\in\mathbb{R}^{d}\}$, where $\eta_{\nabla f^{\star}}=\frac{\nabla f^{\star}(x_{0})\nabla f^{\star}(x_{0})^{\top}}{\lVert\nabla f^{\star}(x_{0})\rVert^{2}}$. Taylor expansion shows $\mathcal{I}_{\text{tangent}}(x_{0})\approx\mathcal{I}(x_{0})$ when restricted to the small neighborhood of $x_{0}$. To generate algorithmic insight, we further define a probability density on $\mathcal{I}_{\text{tangent}}(x_{0})$: we use $P(x^{\prime};x_{0},\sigma)$ to denote the probability density function of $x^{\prime}=x_{0}+\left(I-\eta_{\nabla f^{\star}}\right)z$ where $z$ is sampled from a $d$-dimensional isotropic Gaussian distribution $\mathcal{N}(0,\sigma^{2})$. It is easy to verify that $P(\cdot;x_{0},\sigma)$ is a Gaussian distribution supported on $\mathcal{I}_{\text{tangent}}(x_{0})$ whose mean is $x_{0}$.
@@
Despite its simplicity, Lemma 2.1 demonstrates two important elements of perturbation-based UQ: input perturbation and output variance evaluation. In the following subsections, we introduce designs that implement these elements in the LLM context to characterize the uncertainty.
@@ 2.2 A probabilistic framework via dual random walks
We use $S_{x_{0}}$ to denote a semantic input to an LLM $f$ and $S_{y_{0}}=f(S_{x_{0}})$ one of its possible corresponding outputs. Since $f$ is usually stochastic, the output $S_{y_{0}}=f(S_{x_{0}})$ is also random. Now consider a semantic embedding function $\psi$, which maps input and output samples into a continuous space $\mathcal{X}$: $(x_{0},y_{0})=(\psi(S_{x_{0}}),\psi(S_{y_{0}}))$. Also, assume we have a perturbation algorithm $\mathcal{P}er(S_{x_{0}})$ that generates a finite set of perturbed inputs:
$\mathcal{P}er(S_{x_{0}})=\{S_{x_{0}},S_{x_{1}},\dots,S_{x_{n}}\},$
where the detailed implementation of $\mathcal{P}er(S_{x_{0}})$ will be discussed in Sec. 2.5.
@@
We first introduce two Markov chains $\mathcal{M}_{x}$ and $\mathcal{M}_{y}$. They both contain the same states $\mathcal{S}=\{0,1,2,\cdots,n\}$. Assume a non-negative similarity function $a_{\text{Similarity}}(x,x^{\prime}):\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{0}^{+}$ that quantifies how close are two embeddings in $\mathcal{X}$, then we can define the transition probability matrices of $\text{P}_{x},\text{P}_{y}\in\mathbb{R}^{(n+1)\times(n+1)}$ as:
$\text{P}_{x}[i,j]$ $\triangleq\frac{a_{\text{Similarity}}(x_{i},x_{j})}{\sum_{k=0}^{n}a_{\text{Similarity}}(x_{i},x_{k})},$ (2)
$\text{P}_{y}[i,j]$ $\triangleq\frac{a_{\text{Similarity}}(y_{i},y_{j})}{\sum_{k=0}^{n}a_{\text{Similarity}}(y_{i},y_{k})}.$
The two transition probability matrices characterize two random walks in the space of $n+1$ pairs $\{x_{i},y_{i}\}_{i=0}^{n}$, where the transition probability from $i$ to $j$ is higher if their input or output semantic features are closer. Notice that $a_{\text{Similarity}}$ can be defined in different ways, some of which are explored in our numerical studies.
At its core,  (2) defines stochastic dynamics on the set $\mathcal{S}$, capturing the similarity structures in both the input and output spaces. These dual dynamics uncover meaningful semantic patterns that can be leveraged for UQ. There are various ways to characterize uncertainty by examining the alignment between the two induced graphs (Vishwanathan et al., 2010). In what follows, we construct a framework tailored for discrete input and output spaces to rigorously define uncertainty.
@@ 2.3 Constructing the distributions
The conditional distribution-$P(X|Y)$
Next, we introduce the conditional probability $P(X=x_{i}\mid Y=y_{j})$ as
@@
where notation $[j,i]$ denotes the $(j,i)$-th entry of the matrix $\text{P}_{y}\text{P}_{x}\in\mathbb{R}^{n+1\times n+1}$.
Essentially, (4) constructs the conditional probability using the transition probability from state $i$ to $j$ under the composite dynamics. We design a two-stage process on $\mathcal{S}$. Initially, states are transitioned by $\text{P}_{y}$ that characterizes the output similarity. Then, the states are transitioned by $\text{P}_{x}$ that characterizes the input similarity. Since $\text{P}_{x}$ and $\text{P}_{y}$ are defined on the same state space $\mathcal{S}$ linked by an LLM, (4) establishes a probabilistic bridge linking similarity in the output space to the similarity in the input space. This is shown in Fig. 2.
We explicitly model the conditional distribution of input $X$ rather than output $Y$ for two reasons. First, samples generated by the perturbation algorithm $\mathcal{P}er(S_{x_{0}})$ may deviate from $x_{0}$ in meanings. We thus reweight these samples based on their pairwise similarity in the input space and output space. Second, we observe there exists an inherent asymmetry in semantic variability between inputs and outputs in LLMs: multiple semantically distinct prompts can often elicit very similar outputs, whereas small changes to an input typically result in only modest variation in output. As a result, modeling the conditional distribution of $X$ given $Y$ provides a more stable and reliable basis for inferring the diversity of inputs that are semantically related to an output.
Our goal is to leverage the information from $Y$ to guide the conditional distribution of $X$. This interplay between input and output similarities allows the probability distribution to adapt to the structure of the questions and responses, ensuring that inputs are assigned higher probability when they are supported by multiple semantically similar input-output pairs.
The distributions derived from $P(Y)$ and $P(X|Y)$
Based on defined $P(Y)$ and $P(X|Y)$, we can calculate the joint probability $P(X=x_{i},Y=y_{j})\triangleq P(x_{i}|y_{j})P(y_{j})$, and the marginal probability $P(X=x_{i})\triangleq\sum_{j=0}^{n}P(x_{i}|y_{j})P(y_{j})=\pi^{\text{Uniform}}\text{P}_{y}\text{P}_{y}\text{P}_{x}[i]$. We are also ready to apply Bayes’ theorem to compute the conditional probability $P(Y=y_{j}|X=x_{i})\triangleq P(y_{j}|x_{i})=\frac{P(x_{i}|y_{j})P(y_{j})}{\sum_{k=0}^{n}P(x_{i}|y_{k})P(y_{k})}$.
Building on this framework, researchers can flexibly define uncertainty measures tailored to their needs. One of our main contributions is to offer this general probabilistic foundation based on $P(Y)$ and $P(Y|X)$, which supports a wide range of definitions such as divergence-based metrics (for example, KL divergence and Wasserstein distance), entropy-based measures derived from various distributions, and other probabilistic constructs. As an example, we introduce an entropy-based metric in the following section.

### v2

@@ Abstract
Large language models (LLMs) have transformed natural language processing, but their reliable deployment requires effective uncertainty quantification (UQ). Existing UQ methods are often heuristic and lack a probabilistic interpretation. This paper begins by providing a theoretical justification for the role of perturbations in UQ for LLMs. We then introduce a dual random walk perspective, modeling input–output pairs as two Markov chains with transition probabilities defined by semantic similarity. Building on this, we propose a fully probabilistic framework based on an inverse model, which quantifies uncertainty by evaluating the diversity of the input space conditioned on a given output through systematic perturbations. Within this framework, we define a new uncertainty measure, Inv-Entropy. A key strength of our framework is its flexibility: it supports various definitions of uncertainty measures, embeddings, perturbation strategies, and similarity metrics. We also propose GAAP, a perturbation algorithm based on genetic algorithms, which enhances the diversity of sampled inputs. In addition, we introduce a new evaluation metric, Temperature Sensitivity of Uncertainty (TSU), which directly assesses uncertainty without relying on correctness as a proxy. Extensive experiments demonstrate that Inv-Entropy outperforms existing semantic UQ methods. The code to reproduce the results can be found at https://github.com/UMDataScienceLab/Uncertainty-Quantification-for-LLMs.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success in various natural language processing tasks, such as text generation, question answering, and summarization (Brown2020; Chowdhery2022; Touvron2023). These models have pushed the boundaries of what is achievable in language understanding and generation (Chung2022; OpenAI2023). However, despite their impressive capabilities, a significant challenge remains: LLMs tend to hallucinate, or generate confidently wrong predictions (Maynez2020; Zhang2023). This is a serious concern in applications where reliability is paramount, such as healthcare, autonomous systems, and legal domains, where incorrect outputs can have dire consequences (Ji2023). Addressing these challenges is essential for ensuring the reliable and responsible deployment of LLMs at scale, ultimately unlocking their full potential in real-world applications.
@@
A central step in addressing these limitations is developing effective measures for UQ, enabling LLMs to acknowledge their confidence in a generated output. Existing UQ approaches rely predominantly on heuristic consistency checks. Most commonly, they use the model’s own generation likelihood or perplexity as a proxy for confidence, or they measure dispersion across multiple sampled continuations (e.g. via n‑gram overlap or embedding‑space variance) (mudumbai2024). Such likelihood-based and sampling‑based measures, however, lack a grounded probabilistic justification. Also, token‑level probabilities are known to dramatically under‑estimate uncertainty as models can be “confidently wrong” (jiang-etal-2021-know), and they are often inaccessible in black-box LLMs.
@@
To probe deeper LLM brittleness, recent work has turned to input‑perturbation UQ methods. Variant prompts are created through paraphrasing, adversarial token insertion, or temperature shifts. Output sensitivity is then quantified as an uncertainty signal (gao2024spuq; Tuna2022; Seeböck2019). For instance, gao2024spuq randomly perturb both prompt wording and sampling temperature, then aggregate variation across outputs to flag unstable predictions. Tuna2022 applies adversarial paraphrases to uncover “blind spots” where small semantic-preserving edits trigger large output changes, while Seeböck2019 uses systematic character‑level and word‑level corruptions to map regions of high model vulnerability.
@@
Despite these advances, an ab initio probabilistic framework has not been established, as existing methods mainly rely on heuristic score functions for UQ. To address this, we introduce Inv-Entropy, a fully probabilistic framework built on random-walk theory that offers a new perspective: it learns the statistical connections between LLM inputs and outputs. This is accomplished through structured perturbations, which simultaneously capture input variability and the influence of different inputs on model predictions. Our contributions are summarized as follows:
@@
    We present the first work in UQ for LLMs that adopts a fully probabilistic framework, grounded in random walk theory. This framework is highly flexible, and its probabilistic nature allows for the use of various UQ measures. It is also intrinsically capable of handling black-box models, as it does not rely on token probabilities.
@@
    Theoretically, our work provides a principled foundation for using perturbation-based methods for UQ, justifying the recent momentum behind such approaches.
@@
Figure 1: Toy example highlighting the importance of perturbations. The original question is from TriviaQA (joshi2017triviaqa), and the correct answer is “bras.” The responses are generated by ChatGPT-3.5-Turbo. Input perturbations reveal hidden variability that multiple sampling (i.e., replications) alone fails to capture, as replication alone can be confidently wrong.
@@ 2.1 Why perturb the input?
We begin with a simple illustrative example in Fig. 1 to highlight the importance of input perturbations. In this example, when the input question is simply replicated (as is common in much of the existing literature (lin2024generating)) all generated responses are identical. This misleadingly suggests low uncertainty even though all answers are incorrect, thus providing false confidence. Replication alone is therefore insufficient to capture the underlying variability. In contrast, applying input perturbations yields a diverse set of responses across semantically equivalent prompts, enabling a more faithful characterization of uncertainty. The importance of such perturbations is not only evident from this example but also supported by simplified theoretical argument in what follows.
To build theoretical insight into the importance of perturbations, we leverage a key property specific to language: semantic equivalence. An ideal LLM should respond consistently to semantically equivalent inputs. We use this property, in a simplified setting, to construct input perturbations within equivalence classes, which allows us to expose inconsistencies in model behavior and formally reason about uncertainty. We start from a proof-of-concept example where the target function $f^{\star}$ is a single output function $f^{\star}:\mathbb{R}^{d}\to\mathbb{R}$. We lack access to $f^{\star}$, but have access to an approximation $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}$, such as a pre-trained model. At first sight, it seems hopeless to quantify the alignment between $\hat{f}$ and $f^{\star}$ without knowing $f^{\star}$. However, the semantic equivalence class relative to an input $x_{0}$ provides a set $\mathcal{I}(x_{0})\subseteq\mathbb{R}^{d}$, such that $f^{\star}(x^{\prime})=f^{\star}(x_{0}),\,\forall x^{\prime}\in\mathcal{I}(x_{0})$. The equivalence class provides valuable information for UQ as shown in the subsequent argument and Figure 2.
The shape of $\mathcal{I}(x_{0})$ could be complex for general $f^{\star}$. However, in a small neighborhood of $x_{0}$ where $\nabla f^{\star}(x_{0})\neq 0$, $\mathcal{I}(x_{0})$ should be locally close to the tangent space of $f^{\star}$. More specifically, we define a tangent invariance set as $\mathcal{I}_{\text{tangent}}(x_{0})=\{x_{0}+\left(I-\eta_{\nabla f^{\star}}\right)z;z\in\mathbb{R}^{d}\}$, where $\eta_{\nabla f^{\star}}=\frac{\nabla f^{\star}(x_{0})\nabla f^{\star}(x_{0})^{\top}}{\lVert\nabla f^{\star}(x_{0})\rVert^{2}}$. Intuitively, $I-\eta_{\nabla f^{\star}}$ acts as the orthogonal projector onto the subspace orthogonal to $\nabla f^{\star}(x_{0})$. Taylor expansion shows $\mathcal{I}_{\text{tangent}}(x_{0})\approx\mathcal{I}(x_{0})$ when restricted to the small neighborhood of $x_{0}$. To generate algorithmic insight, we further define a probability density on $\mathcal{I}_{\text{tangent}}(x_{0})$: we use $P(x^{\prime};x_{0},\sigma)$ to denote the probability density function of $x^{\prime}=x_{0}+\left(I-\eta_{\nabla f^{\star}}\right)z$ where $z$ is sampled from a $d$-dimensional isotropic Gaussian distribution $\mathcal{N}(0,\sigma^{2})$. It is easy to verify that $P(\cdot;x_{0},\sigma)$ is a Gaussian distribution supported on $\mathcal{I}_{\text{tangent}}(x_{0})$ whose mean is $x_{0}$.
@@
Figure 2: Left: Conceptual illustration of level sets of the ground truth $f^{\star}$. We perturb the input $x_{0}$ to $x_{1},x_{2},\ldots$ along a schematic isocontour such that $f^{\star}(x_{0})=f^{\star}(x_{1})=\cdots$. Right: Conceptual illustration of level sets of the model $\hat{f}$. The deviations of $\hat{f}(x_{i})$ for $i\geq 1$ from $\hat{f}(x_{0})$ reflect the model’s uncertainty around $x_{0}$.
@@
Despite its simplicity, Lemma 2.1 demonstrates two important elements of perturbation-based UQ: input perturbation and output variance evaluation. In what follows, we introduce concrete designs that implement these elements in the LLM context to effectively characterize uncertainty.
@@ 2.2 A probabilistic framework via dual random walks
We use $S_{x_{0}}$ to denote a semantic input to a LLM $f$, and $S_{y_{0}}\leftarrow f(S_{x_{0}})$ to denote one of its possible corresponding outputs. Because $f$ is typically stochastic, the output $S_{y_{0}}$ is a random variable. Next, consider a semantic embedding function $\psi$ that maps both inputs and outputs into a continuous space $\mathcal{X}$, such that $(x_{0},y_{0})=(\psi(S_{x_{0}}),\psi(S_{y_{0}})).$ We also assume a perturbation algorithm $\mathcal{P}\mathrm{er}(S_{x_{0}})$ that produces a finite set of perturbed inputs:
$\mathcal{P}\mathrm{er}(S_{x_{0}})=\{S_{x_{0}},S_{x_{1}},\dots,S_{x_{n}}\,\},$
whose detailed implementation will be described in Sec. 2.6.
@@
We define two Markov chains, $\mathcal{M}_{x}$ and $\mathcal{M}_{y}$, over the same state set $\mathcal{S}=\{0,1,\dots,n\}$, where each state $i$ corresponds to a perturbed instance $S_{x_{i}}$. The chains differ in their transition dynamics: $\mathcal{M}_{x}$ is built from similarities among input embeddings $X_{n}$, and $\mathcal{M}_{y}$ from similarities among output embeddings $Y_{n}$. Let $a_{\text{Similarity}}(x,x^{\prime}):\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{0}^{+}$ denote a non-negative similarity function that measures the closeness between embeddings in $\mathcal{X}$. Using this function, we define the transition matrices $\text{P}_{x},\text{P}_{y}\in\mathbb{R}^{(n+1)\times(n+1)}$ for $\mathcal{M}_{x}$ and $\mathcal{M}_{y}$ elementwise as:
$\text{P}_{x}[i,j]$ $\triangleq\frac{a_{\text{Similarity}}(x_{i},x_{j})}{\sum_{k=0}^{n}a_{\text{Similarity}}(x_{i},x_{k})},\quad\text{P}_{y}[i,j]$ $\triangleq\frac{a_{\text{Similarity}}(y_{i},y_{j})}{\sum_{k=0}^{n}a_{\text{Similarity}}(y_{i},y_{k})}.$ (2)
The two transition probability matrices characterize two random walks in the space of $n+1$ pairs $\{x_{i},y_{i}\}_{i=0}^{n}$, where the transition probability from $i$ to $j$ is higher if their input or output semantic features are closer. Notice that a variety of well-established methods have been proposed for $a_{\text{Similarity}}$, some of which are deployed in our numerical studies (see Sec. 3).
At its core,  (2) defines stochastic dynamics on the set $\mathcal{S}$, capturing the similarity structures in both the input and output spaces. These dual dynamics uncover meaningful semantic patterns that can be leveraged for UQ. There are various ways to characterize uncertainty by examining the alignment between the two induced graphs (graphkernels). In what follows, we construct a framework tailored for discrete input and output spaces to rigorously define uncertainty.
@@ 2.3 Constructing the distributions
The conditional distribution
Now, we introduce the conditional probability $P(X=x_{i}\mid Y=y_{j})$
@@
where notation $[j,i]$ denotes the $(j,i)$-th entry of the matrix $\text{P}_{y}\text{P}_{x}\in\mathbb{R}^{(n+1)\times(n+1)}$.
Figure 3: Random-walk transitions underlying $P(X\mid Y)=\text{P}_{y}\text{P}_{x}$. Highlighted blue paths show two representative transitions (one through $k$ and one through $n$), each following $y_{j}\xrightarrow{\text{P}_{y}}y_{k}\xrightarrow{\text{LLM}}x_{k}\xrightarrow{\text{P}_{x}}x_{i}$.
Essentially,  (4) defines the conditional probability through composite transition dynamics on the state space $\mathcal{S}$. We design a two-stage random walk: first, states transition under $\text{P}_{y}$, capturing output similarities; then, they transition under $\text{P}_{x}$, capturing input similarities. Since both $\text{P}_{x}$ and $\text{P}_{y}$ operate on the same state space $\mathcal{S}$ linked by the LLM,  (4) establishes a probabilistic bridge connecting similarity structures in the output and input spaces (see Fig. 3).
We explicitly model the conditional distribution of inputs $X$ rather than outputs $Y$ for two main reasons. First, the perturbed samples generated by $\mathcal{P}\mathrm{er}(S_{x_{0}})$ may differ semantically from the original input. Reweighting these samples using both input- and output-space similarities ensures that semantically consistent perturbations are emphasized while spurious ones are down-weighted. Second, LLMs exhibit an inherent semantic asymmetry between inputs and outputs: many distinct prompts can lead to similar responses, whereas small input changes typically cause only small to modest output variations. Modeling $P(X\mid Y)$ therefore provides a more stable and informative view of uncertainty by capturing the diversity of possible inputs that could relate to a given output.
Our goal is to use the information encoded in $Y$ to guide the conditional distribution of $X$. This coupling between input and output similarities allows the model to assign higher probability to inputs supported by multiple semantically consistent input-output pairs, yielding a more faithful representation of uncertainty. Based on the defined $P(Y)$ and $P(X\mid Y)$, we can then derive $P(X)\triangleq\pi^{\text{Uniform}}\text{P}_{y}\text{P}_{x}$ and $P(Y\mid X)$ via Bayes’ theorem. Such formulation provides a flexible foundation for defining diverse uncertainty measures, including divergence-based metrics (e.g., KL or Wasserstein distances), entropy-based quantities, and other probabilistic constructs. In the next section, we introduce an entropy-based metric as a concrete example.